# Data Efficient Adaptation in Large Language Models via Continuous Low-Rank Fine-Tuning

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have emphasized the critical role of fine-tuning (FT) techniques in adapting LLMs to specific tasks, especially when retraining from scratch is computationally infeasible. Fine-tuning enables LLMs to leverage task- or domain-specific data, producing models that more effectively meet the requirements of targeted applications. However, conventional FT approaches often suffer from catastrophic forgetting and suboptimal data efficiency, limiting their real-world applicability. To address these challenges, this paper proposes DEAL, a novel framework that integrates Low-Rank Adaptation (LoRA) with a continuous fine-tuning strategy. By incorporating knowledge retention and adaptive parameter update modules, the framework mitigates the limitations of existing FT methods while maintaining efficiency in privacy-preserving settings. Experiments on 15 diverse datasets show that DEAL consistently outperforms baseline methods, yielding substantial gains in task accuracy and resource efficiency. These findings demonstrate the potential of our approach to advance continual adaptation in LLMs by enhancing task performance while improving resource efficiency. The source code is publicly available at https://github.com/zzm-black/DEAL-Continuous-Low-Rank-Fine-Tuning.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has catalyzed transformative advances in Natural Language Processing (NLP), enabling breakthroughs across healthcare, education, web technologies, and other domains Luo et al. (2022); Yang et al. (2023). However, training and utilizing these models to evolve to real-world demands remains a critical challenge. Direct fine-tuning of billion-scale-parameter models incurs prohibitive computational costs, creating accessibility barriers for resource-constrained researchers and institutions. Even for small and medium-sized enterprises, it is also difficult to independently implement pre-training of these models. In such scenarios, Parameter-Efficient Fine-Tuning (PEFT) methods—particularly Low-Rank Adaptation (LoRA)—have emerged as practical solutions that leverage pre-trained models by “standing on the shoulders of giants” Hu et al. (2021). By applying low-rank matrix decompositions, LoRA reduces the number of trainable parameters by over 90% while maintaining baseline performance. By selectively updating task-specific subspaces, LoRA enables targeted knowledge integration in non-stationary environments—a capability aligned with the core objectives of continual learning.
@@ 3.2 Wavelet Kernel-based Knowledge Retention
Here, $\boldsymbol{X}$ denotes the core feature matrix, capturing the intrinsic low-rank structure that encodes task-relevant semantics. According to the Eckart–-Young–-Mirsky theorem  Eckart and Young (1936), the best low-rank approximation of a matrix in the Frobenius norm sense is achieved via truncation of its singular value decomposition (SVD). This motivates our use of truncated SVD to estimate $\boldsymbol{X}$ from the observed matrix $\boldsymbol{Y}$, with the goal of recovering task-relevant features from its singular representation.
@@
where $\boldsymbol{u}_{k}$ and $\boldsymbol{v}_{k}$ is the left and right singular vectors of $\boldsymbol{Y}$, $\sigma_{k}$ is the $k$-th largest singular value of $\boldsymbol{Y}$ ($\sigma_{1}>\sigma_{2}>\cdots>\sigma_{r}$). In traditional signal analysis algorithms, large singular values represent low-frequency data distribution and macro trends, and small singular values represent high-frequency disturbances Liao and Xu (2017). However, since $\sigma^{2}$ in the above formula is unknown, we will define a series of wavelet functions at different scales for feature filtering. Here, we use the heat kernel as a low-pass filter:
@@ 4 Experiments
(iii) Heterogeneous multi-task learning: We evaluate on the 15-task benchmark proposed by Asai et al. (2023), which spans text classification (AG News, DBpedia, Yahoo, Amazon, Yelp), GLUE tasks (MNLI, QQP, RTE, SST-2) Wang et al. (2018), SuperGLUE tasks (WiC, CB, COPA, MultiRC, BoolQ) Wang et al. (2019), and IMDB Maas et al. (2011). This benchmark tests DEAL’s scalability and robustness across heterogeneous tasks and long task sequences. Full details on dataset preprocessing, task ordering, and prompt construction are provided in Appendix B.1D.
@@ B.1 Datasets
Table 3 summarizes all datasets used across the continual learning benchmarks. These datasets were also employed in O-LoRA [Wang et al., 2023a], where each is framed as a classification task using a unified instruction-based text-to-text format.
@@ B.2 Baselines
  • O-LoRA Wang et al. [2023a]: A recent method that allocates task-specific adapters to orthogonal subspaces, reducing parameter interference.
@@ 3 Yahoo $\rightarrow$ Amazon $\rightarrow$ AG News $\rightarrow$ DBpedia
Appendix D Instruction Prompts
We adopt the task-specific instruction prompts introduced in O-LoRA [Wang et al., 2023a], as summarized in Table 6.
Table 6: Instruction prompts provided to the model for each task.
Task Prompt
Appendix E Case Study Examples
@@ Appendix G Comparison with Additional Continual Learning Methods
  • Replay: Rehearses past samples to prevent forgetting by storing and replaying them during training Lopez-Paz and Ranzato [2017].
  • EWC: Mitigates forgetting by penalizing changes to important parameters estimated via Fisher information Kirkpatrick et al. [2017].
  • LwF: Preserves knowledge of past tasks through knowledge distillation without storing old data Li and Hoiem [2017].
  • L2P: Introduces learnable prompts for continual learning to guide pre-trained models effectively Wang et al. [2022].
  • LFPT5: Proposes a unified framework for lifelong few-shot learning based on prompt tuning of T5 Qin and Joty [2021].
  • ProgPrompt: Designs progressive prompts to adapt large language models for continual learning Razdaibiedina et al. [2023].
  • LB-CL: Achieves parameter-efficient continual learning by learning more but disturbing less Qiao and Mahdavi [2024].

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) have emphasized the critical role of fine-tuning (FT) techniques in adapting LLMs to specific tasks, especially when retraining from scratch is computationally infeasible. Fine-tuning enables LLMs to leverage task- or domain-specific data, producing models that more effectively meet the requirements of targeted applications. However, conventional FT approaches often suffer from catastrophic forgetting and suboptimal data efficiency, limiting their real-world applicability. To address these challenges, this paper proposes DEAL, a novel framework that integrates Low-Rank Adaptation (LoRA) with a continuous fine-tuning strategy. By incorporating knowledge retention and adaptive parameter update modules, the framework mitigates the limitations of existing FT methods while maintaining efficiency. Experiments on 15 diverse datasets show that DEAL consistently outperforms baseline methods, yielding substantial gains in task accuracy and resource efficiency. These findings demonstrate the potential of our approach to advance continual adaptation in LLMs by enhancing task performance while improving resource efficiency. The source code is publicly available at https://github.com/zzm-black/DEAL-Continuous-Low-Rank-Fine-Tuning.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has catalyzed transformative advances in Natural Language Processing (NLP), enabling breakthroughs across healthcare, education, web technologies, and other domains Luo et al. (2022); Yang et al. (2023); Fu et al. (2025). However, training and utilizing these models to evolve to real-world demands remains a critical challenge. Direct fine-tuning of billion-scale-parameter models incurs prohibitive computational costs, creating accessibility barriers for resource-constrained researchers and institutions. Even for small and medium-sized enterprises, it is also difficult to independently implement pre-training of these models. In such scenarios, Parameter-Efficient Fine-Tuning (PEFT) methods—particularly Low-Rank Adaptation (LoRA)—have emerged as practical solutions that leverage pre-trained models by “standing on the shoulders of giants” Hu et al. (2021). By applying low-rank matrix decompositions, LoRA reduces the number of trainable parameters by over 90% while maintaining baseline performance. By selectively updating task-specific subspaces, LoRA enables targeted knowledge integration in non-stationary environments—a capability aligned with the core objectives of continual learning.
@@ 3.2 Wavelet Kernel-based Knowledge Retention
Here, $\boldsymbol{X}$ denotes the core feature matrix, capturing the intrinsic low-rank structure that encodes task-relevant semantics. According to the Eckart–-Young–-Mirsky theorem  Eckart, Young (1936), the best low-rank approximation of a matrix in the Frobenius norm sense is achieved via truncation of its singular value decomposition (SVD). This motivates our use of truncated SVD to estimate $\boldsymbol{X}$ from the observed matrix $\boldsymbol{Y}$, with the goal of recovering task-relevant features from its singular representation.
@@
where $\boldsymbol{u}_{k}$ and $\boldsymbol{v}_{k}$ is the left and right singular vectors of $\boldsymbol{Y}$, $\sigma_{k}$ is the $k$-th largest singular value of $\boldsymbol{Y}$ ($\sigma_{1}>\sigma_{2}>\cdots>\sigma_{r}$). In traditional signal analysis algorithms, large singular values represent low-frequency data distribution and macro trends, and small singular values represent high-frequency disturbances Liao, Xu (2017). However, since $\sigma^{2}$ in the above formula is unknown, we will define a series of wavelet functions at different scales for feature filtering. Here, we use the heat kernel as a low-pass filter:
@@ 4 Experiments
(iii) Heterogeneous multi-task learning: We evaluate on the 15-task benchmark proposed by Asai et al. (2023), which spans text classification (AG News, DBpedia, Yahoo, Amazon, Yelp), GLUE tasks (MNLI, QQP, RTE, SST-2) Wang et al. (2018), SuperGLUE tasks (WiC, CB, COPA, MultiRC, BoolQ) Wang et al. (2019), and IMDB Maas et al. (2011). This benchmark tests DEAL’s scalability and robustness across heterogeneous tasks and long task sequences. Full details on dataset preprocessing, task ordering, and prompt construction are provided in Appendix B.1E.
@@ 7 Conclusion
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (No. 62502404), the Research Impact Fund of the Hong Kong Research Grants Council (No. R1015-23), the Collaborative Research Fund of the Research Grants Council (No. C1043-24GF), the General Research Fund of the Research Grants Council (No. 11218325), and the Institute of Digital Medicine at City University of Hong Kong (No. 9229503). Additional support was provided by Huawei (Huawei Innovation Research Program), Tencent (CCF–Tencent Open Fund; Tencent Rhino-Bird Focused Research Program), Alibaba (CCF–Alimama Tech Kangaroo Fund No. 2024002), Ant Group (CCF–Ant Research Fund), Didi (CCF–Didi Gaia Scholars Research Fund), Kuaishou, and ByteDance.
@@ B.1 Datasets
Table 3 summarizes all datasets used across the continual learning benchmarks. These datasets were also employed in O-LoRA (Wang et al., 2023a), where each is framed as a classification task using a unified instruction-based text-to-text format.
@@ B.2 Baselines
  • O-LoRA Wang et al. (2023a): A recent method that allocates task-specific adapters to orthogonal subspaces, reducing parameter interference.
@@ 3 Yahoo $\rightarrow$ Amazon $\rightarrow$ AG News $\rightarrow$ DBpedia
Appendix D Case Study Examples
@@
Appendix E Instruction Prompts
We adopt the task-specific instruction prompts introduced in O-LoRA (Wang et al., 2023a), as summarized in Table 6.
Table 6: Instruction prompts provided to the model for each task.
Task Prompt
@@ Appendix G Comparison with Additional Continual Learning Methods
  • Replay: Rehearses past samples to prevent forgetting by storing and replaying them during training Lopez-Paz, Ranzato (2017).
  • EWC: Mitigates forgetting by penalizing changes to important parameters estimated via Fisher information Kirkpatrick et al. (2017).
  • LwF: Preserves knowledge of past tasks through knowledge distillation without storing old data Li, Hoiem (2017).
  • L2P: Introduces learnable prompts for continual learning to guide pre-trained models effectively Wang et al. (2022).
  • LFPT5: Proposes a unified framework for lifelong few-shot learning based on prompt tuning of T5 Qin, Joty (2021).
  • ProgPrompt: Designs progressive prompts to adapt large language models for continual learning Razdaibiedina et al. (2023).
  • LB-CL: Achieves parameter-efficient continual learning by learning more but disturbing less Qiao, Mahdavi (2024).